# Context-Aware Deep Learning for Robust Channel Extrapolation in Fluid Antenna Systems

### v2

@@ III-B5 Loss Function
where $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\mathrm{FFT}}$ are the mean-square-error (MSE) loss and the FFT loss attributed to spatial amplitude perturbation, respectively, calculated as
@@ IV Numerical Results
The performance is evaluated through the normalized MSE (NMSE), defined as
@@
(a) $W_{s}=2~{}{\rm cm}\times 4~{}{\rm cm}$
(b) $W_{s}=8~{}{\rm cm}\times 16~{}{\rm cm}$
Figure 2: NMSE versus the number of observed ports with different FAS sizes.
Figs. 2(a) and 2(b) illustrate the NMSE curves of different models under various SNR conditions, as a function of the observed port ratio, for FAS sizes of $2~{}{\rm cm}\times 4~{}{\rm cm}$ and $8~{}{\rm cm}\times 16~{}{\rm cm}$, respectively. We reproduce the AGMAE in [25] as a baseline for comparison. The simplified version of CANet without spatial amplitude perturbation and spectral consistency loss is denoted as ‘CANet-B’, while the full CANet model includes all the proposed components. In the figures, curves with the same line style represent the same SNR setting.
The results demonstrate that as the ratio of observed ports increases, the NMSE of all the models decreases considerably, showing that richer observed information leads to better reconstruction performance. Under extremely low observation rates (e.g., $5\%$), AGMAE slightly outperforms CANet-B, suggesting that its structure has certain advantages under highly sparse inputs. But in most observation rates, CANet-B consistently surpasses AGMAE. Furthermore, with the incorporation of amplitude spectrum perturbation and spectral consistency loss, the overall performance of CANet is further improved. For example, at an observation rate of $20\%$ with the $2~{}{\rm cm}\times 4~{}{\rm cm}$ FAS, CANet achieves up to a $3~{}{\rm dB}$ NMSE gain over CANet-B, demonstrating stronger noise robustness and generalization ability. Fig. 2(b) shows a similar trend under a larger FAS size, validating the adaptability and effectiveness of the proposed method across different physical FAS configurations.

### v4

@@ III-B5 Loss Function
where $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\mathrm{FFT}}$ are the mean squared error (MSE) loss and the FFT loss attributed to spatial amplitude perturbation, respectively, calculated as
@@ IV Numerical Results
IV-A Simulation settings
TABLE I: Architecture of the proposed CANet.
Layers Operator Out Ch. K(S) Act.
1 CAB $32{\times}16$ 64 $5(1)$ LReLU
2 CAB $32{\times}16$ 128 $3(2)$ LReLU
3 CAB $16{\times}8$ 128 $3(1)$ LReLU
4–5 CAB $\times 2$ $16{\times}8$ 256 $3(1)$ LReLU
6 CSCA $16{\times}8$ 256
7–16 ConvNeXt v2 $\times 10$ $16{\times}8$ 256 $7(3)$ GELU
17 BilinearUp $32{\times}16$ 256
18–19 CAB $\times 2$ $32{\times}16$ 128 $3(1)$ LReLU
20–21 CAB $\times 2$ $32{\times}16$ 64 $3(1)$ LReLU
22 Conv $32{\times}16$ 16 $7(1)$
Note: K: kernel size; S: stride; LReLU: LeakyReLU; GELU: Gaussian Error Linear Unit.
@@ 22 Conv $32{\times}16$ 16 $7(1)$
The performance is evaluated through the NMSE, defined as
@@
(a) 3.4GHz
(b) 28GHz
Figure 2: NMSE versus the number of observed ports with different operating frequencies. CANet consistently outperforms benchmark models across all observation ratios, demonstrating superior robustness in reconstructing high-resolution CSI from sparse observations.
IV-B Performance analysis
Fig. 2 illustrates the NMSE performance of different models as a function of the observation ratio for a $2~\mathrm{cm}\times 4~\mathrm{cm}$ FAS array under multiple SNR conditions, where Figs. 2(a) and 2(b) correspond to the carrier frequencies of $3.4$ GHz and $28$ GHz, respectively. We reproduce AGMAE in [37] as a representative deep-learning-based baseline tailored for FAS scenarios, and additionally include CT[39] as a comparison model, which was originally designed for channel extrapolation in conventional multiple-input multiple-output (MIMO) systems, to evaluate the applicability of generic extrapolation models to FAS settings. In the figure, the same line style is used to indicate the same SNR condition.
For both carrier frequencies, the NMSE of all models decreases monotonically with the observation ratio, indicating that richer port observations improve CSI reconstruction accuracy. However, under sparse observations, especially in low-SNR conditions, CT exhibits limited performance gains, suggesting that directly reusing MIMO-oriented extrapolation structures is insufficient to capture the fine-grained port correlations inherent in FAS. In contrast, AGMAE demonstrates more stable extrapolation performance across both frequency bands, while the proposed CANet consistently achieves the lowest NMSE under all SNR and observation ratio settings, highlighting its adaptability to different carrier frequencies.
To further evaluate the impact of channel extrapolation accuracy on communication performance, the outage probability is evaluated based on the signal model and port selection criteria described in [27]. Fig. 3 presents the outage probability (OP) as a function of the SINR threshold under a fixed observation ratio of $10\%$ for different methods. Here, Baseline denotes the case where transmission is performed using full but noisy CSI, while Reference corresponds to direct transmission based only on the observable port CSI. Figs. 3 correspond to observation input SNRs of 0 dB and $20$ dB, respectively.
Figure 3: Outage probability versus SINR threshold at a fixed observation ratio of $10\%$ under 0 dB and 20 dB observation SNR. CANet achieves the lowest outage probability among extrapolation schemes, effectively mitigating noise to approach the performance of direct transmission.
Under both SNR settings, CANet achieves the lowest OP over the entire SINR threshold range. Although the Baseline exploits full-port observations, its performance is still limited by measurement noise; in contrast, CANet effectively suppresses observation noise during extrapolation by learning the spatial correlation and statistical structure of FAS channels, thereby achieving improved communication reliability. As the observation input SNR increases to $20$ dB, the performance gap among different methods narrows, while CANet consistently maintains a clear advantage, demonstrating its robustness under varying noise conditions.
TABLE II: Ablation Study of Physics-Informed Modules ($2~\mathrm{cm}\times 4~\mathrm{cm}$ FAS, 20% Observation Ratio)
ID Modules NMSE (dB)
Perturb. FFT Loss 0 dB 10 dB 20 dB
1 6.22-6.22 15.88-15.88 18.86-18.86
2 6.51-6.51 16.42-16.42 19.45-19.45
3 7.35-7.35 17.46-17.46 20.47-20.47
4 $\textbf{8.50}$ $\textbf{18.85}$ $\textbf{21.83}$
[…]
IV-C Ablation experiments
To verify the effectiveness of the proposed FAS-motivated components, an ablation study is conducted as shown in Table II. Compared to the basic model (ID 1), the independent inclusion of the FFT loss (ID 2) and the spatial amplitude perturbation strategy (ID 3) both lead to significant NMSE reductions across various SNR levels. Specifically, the full CANet (ID 4) achieves the best performance, with an additional gain of approximately 3 dB at 20 dB SNR compared to ID 1. These results confirm that explicitly enforcing frequency-domain consistency and simulating spatial uncertainty effectively guide the network to capture the intrinsic propagation physics rather than overfitting to pixel-wise noise.